David’s utility function for good X and Y is given by 𝑈(𝑥, 𝑦) = 𝑥

2𝑦

3

. Where 𝑃𝑥 , 𝑃𝑦 and 

I are the price of good X, price of good Y and consumer income respectively. 

a. Write the budget Constraint of the consumer.

b. Drive the demand functions for good X and Y

c. What combination of X and Y maximizes the consumer’s at I=100, 𝑃𝑥 =

4, 𝑎𝑛𝑑 𝑃𝑦 = 5

d. Calculate the marginal rate of substitution between X and Y at equilibrium and 

interpret your results.

e. Suppose David faces a new utility function such as 𝑈(𝑥, 𝑦) = 25𝑥

0.25𝑦

0.75 for 

consuming commodity X and Y, redo question b. and c. 

f. Suppose David faces another new utility function such as 𝑈(𝑥, 𝑦) = 15𝑥𝑦

2

for 

consuming commodity X and Y, redo question b. and c. 


1
Expert's answer
2021-06-03T13:04:49-0400

A

The budget constraint is 

"Xp_x+ Y p_y = I"


B.

When px = 4 , py = 5 and I = 100, 

the budget constraint becomes, 4x + 5y = 100

Hence, we need to maximize U(x, y) = x2 y3 subject to 4x + 5y = 100

We know, for two goods, utility maximizing condition is when the ratio of marginal utilities is equal to the price ratio i.e. 


"\frac{MUx }{ MUy} = \frac{px }{py}"


MUx is obtained by partial differentiation of U(x, y) wrt x

Therefore, "MUx = 2 x y^3"

Similarly, "MUy = 3 x^2 y^2"


Hence,"\frac{MUx }{ MUy }= \frac{2 x y^3 }{3 x^2 y^2}= \frac{2}{3}(\frac{y}{x}) = \frac{2y }{ 3x}"


C.

Therefore, utility maximizing condition is:

"\frac{2y}{ 3x} = \frac{p_x }{p_y

}= \frac{4 }{ 5}"

i.e. "5y = 6x"

From the budget equation, replacing 6x in place of 5y we get,

 "4x + 6x = 100\\
 10x = 100\\

 x = 10\\

y = 12"

Hence, the combination of X and Y that maximizes consumer's utility is x = 10 units and y = 12 units.


D.

marginal rate of substitution between X and Y at equilibrium is actually the value of (MUx / MUy) at equilibrium. Obviously at equilibrium this is same as price ratio which is (4/5) or 0.8

MRS is the slope of the indifference curve which, at equilibrium is going to be same as the slope of the budget line which is the price ration. The reason behind this is that, at equilibrium, budget line is tangent to the IC. 

Hence, the value of MRS at equilibrium is going to be 0.8


E.


"MU_X=\frac{\delta U}{\delta X}=25(0.2)X^{0.2-1}Y^{0.75}"


"MU_Y=\frac{\delta U}{\delta Y}=25X^{0.2}(0.75)Y^{0.75-1}"


"MRS=\frac{MU_X}{MU_Y}=\frac{0.2X^{-0.8}Y^{0.75}}{0.75X^{0.2}Y^{-0.25}}"



"MRS=\frac{P_X}{P_Y}"


or "MRS=\frac{4Y}{5X}" gives,


"\frac{4Y}{5X}=\frac{P_X}{P_Y}"


"Y=\frac{5X.P_X}{4P_Y}"


"X=\frac{4Y.P_Y}{5P_X}"


These are the generalized demand functions for good X and Y.


Given:

I=100, Px=4,and Py=5

So the budget constraint becomes 4X+5Y=100

 

Substituting these values in Ywe get


"Y=\frac{5X}{4}\times\frac{4}{5}"

or

Y=X

thus

"4X+5X=100\\X=Y=11.11"



F.

"MU_X=\frac{\delta U}{\delta X}=15 Y^2"

"MU_Y=\frac{\delta U}{\delta Y}=30XY"


"MRS=\frac{MU_X}{MU_Y}=\frac{15Y^2}{30XY}"


"\frac{Y}{2X}=\frac{P_X}{P_Y}=\frac{4}{5}"


"8X=5Y\\4X+8X=100\\12X=100\\X=8.333\\Y=13.333"

Hence, the combination of X and Y that maximizes consumer's utility is x = 8 units and y = 13 units.